# Reservoir-Computing Model for Mapping and Forecasting Neuronal Interactions from Electrophysiological Data

### 23 Jan 2024 v2

@@ Abstract
Electrophysiological nature of neuronal networks allows to reveal various interactions between different cell units at a very short time-scales. One of the many challenges in analyzing these signals is to retrieve the morphology and functionality of a given network. In this work we developed a computational model, based on Reservoir Computing Network (RCN) architecture, which decodes the spatio-temporal data from electro-physiological measurements of neuronal cultures and reconstructs the network structure on a macroscopic domain, representing the connectivity between neuronal units. We demonstrate that the model can predict the connectivity map of the network with higher accuracy than the common methods such as Cross-Correlation and Transfer-Entropy. In addition, we experimentally demonstrate the ability of the model to predict a network response to a specific input, such as localized stimulus.
@@ 1 Introduction
Electrophysiological study in neuroscience provides a wide-vision of the interplay between cells of different types at different scales [1]. Such studies vary from investigating the function of a single cell up to studying the dynamics of complex systems consisting of a large number of cells [2], in the pursuit of obtaining a comprehensive picture of the brain activity. As the complexity of the biological system increases, it becomes more and more challenging to analyze or model the behavior in such systems. Numerous models are designed to picture the dynamics behind neuronal activity, starting from single cell models (e.g., Hodgkin–Huxley model [3]) up to models of large populations [4, 5]. Various methods focus on the biophysical properties of the cells (e.g. membrane voltage), while others focus on the point-process of information propagation (e.g., spike trains). Some approaches use experimental observations to adapt a model which will be a computational counterpart to the biological system [6, 7, 8]. Such methods use Machine- or Deep- Learning techniques to train a given model to construct the desired outcome. While for some research questions such approach could be very inefficient and/or computationally expensive, for others it can provide a practical solution to construct a computational tool for various applications.
@@
We propose in this work a simplified approach for interpreting electrophysiological signals from complex networks, which retrieves the connectivity between the different sampled regions in the network and learns the dynamic interactions between them. The model is based on Reservoir Computer Network (RCN) [9], exploiting the information emerged by sampling electrophysiological signals from a complex neural circuitry. The complexity of these circuits cannot be easily understood from a standard measurement analysis, and hence they are modeled as nonlinear networks with inner random connections. We show that using this model we obtain the connectivity map (CM) with higher accuracy than the most common methods, such as Cross-Correlagram (CC) [10, 11] and Transfer-Entropy (TE) [12]. We also demonstrate the capacity of the model to predict the spatio-temproal response of a given network to a specific input. The model’s predictions underwent evaluation through a combination of experimental measurements and numerical simulations. Experimental settings involved microelectrode array (MEA) recordings of in-vitro mice cortical neurons. Additionally, numerical simulations were conducted using the NEST simulator [13].
2 Results
2.1 Reservoir Computing Model for Electrophysiological Signals Decoding
@@
The Reservoir Computing (RC) model is designed to unveil multichannel electrophysiological data obtained from biological neuronal networks. This model employs a recurrent spiking neural network with a leaky integrator, undergoing training using experimentally recorded multichannel electrophysiological signals. The primary objective of the RC model is to decode these signals, ultimately reconstructing a network structure within the measured domain. In the context of the measurement, where each channel captures signals from a neuronal microcircuit (see Fig. 0(a)), we adopt a representation wherein each such microcircuit or population is treated as a node in the network that is to be reconstructed. This approach enables the RC model to unveil the intricate connectivity and relationships within the biological neuronal networks.
@@
We consider discrete time domain denoted by $n=1,2,3...$ corresponding to real time units ($t$) by the relation:
$t=n\cdot t_{int}$ (1)
where $t_{int}$ is defined as the integration time of the network, representing a distinctive time constant that characterizes the duration for a neuronal circuit to process incoming information (Fig. 0(b)). Assuming that the nodes of the network carry information encoded in the firing rate of its population, let us define the network state by a vector $\mathbf{y}[n]\in\mathbb{R}^{N_{ch}\times 1}$, which represents the instantaneous spike rate (ISR) observation of each of the network channels at time step $n$ ($N_{ch}$ is the number of measurement channels). For the prediction of the network next state $\mathbf{y}[n+1]$ as a function of the current and past states $\mathbf{y}[n],\mathbf{y}[n-1],\mathbf{y}[n-2]\ldots$, we designed the artificial neural network (ANN) depicted in Fig. 0(c).
The following equations describe the modeled data processing from state $\tilde{\mathbf{y}}[n]$ to $\tilde{\mathbf{y}}[n+1]$ (the tilde sign indicates predicted states) implementing a reservoir computing network (RCN) approach:
@@
  • $\mathbf{x}_{in},\mathbf{x}\in\mathbb{R}^{N_{res}\times 1}$ are the reservoir input and actual state respectively ($N_{res}$ is the reservoir dimension).
  • $\mathcal{W}_{in}\in\mathbb{R}^{N_{res}\times N_{ch}},\hat{\mathcal{S}},\mathcal{W}_{res}\in\mathbb{R}^{N_{res}\times N_{res}},\mathcal{W}_{out}\in\mathbb{R}^{N_{ch}\times N_{res}}$ are input, synaptic, reservoir and output matrices respectively.
  • $10<\alpha<$ 1 is the memory strength.
  • $\mathbf{b}\in\mathbb{R}^{N_{ch}\times 1}$ is the bias vector.
@@
To train the described ANN, we used experimental recordings and simulated activity representing multichannel ISR sequences as training data. Taking advantage of the RCN paradigm which states that the nonlinear reservoir unit within the ANN inherently possesses dynamic richness, allows to apply the training only on the output layer of the ANN (Eq. (2c)), leveraging the intricate dynamics embedded in the reservoir unit (Eqs. (2a),(2b)). We consider a linear output layer and thereby applying training through a linear regression approach.
Figure 1: Description of the RC model. (a) A microscope image of microelectrode (diameter of 30 $\mu$m) surrounded by cultured neurons. This image illustrates how each electrophysiological measurement site samples signals from a complex circuit, whose morphology is not resolvable by simple methods. (b) An illustrative image of how the signals are intercepted by each electrode (yellow circles) from the biological neuronal network in the background. The spike trains collected at each electrode are then forming integrated signals which propagate within the network. (c) The artificial neural network (ANN) design of the RC model. Consists of three layers (input, reservoir, output) and a recurrent branch. It describes how the state of the network $\mathbf{y}$ (described by $N_{ch}$ samples of the measured signals) is processed from time step $n$ to time step $n+1$. (d) Enhanced visual representation illustrating the interconnection between the computational circuitry of the RC model and the intricate circuitry of biological signals. (e) A block diagram illustrating the sequential computational processes, including training, validation, and testing, conducted in this study.
Let us consider the training data matrix $\mathbf{Y}\in\mathbb{R}^{N_{ch}\times N_{t}}$ which governs $N_{t}$ time steps of the observed network state, i.e. $\mathbf{Y}=\Bigl{[}\mathbf{y}[1],\mathbf{y}[2]\ldots\mathbf{y}[N_{t}]\Bigr{]}$. The task of the training is to optimize output layer’s matrix weights $\mathcal{W}_{out}$ and biases $\mathbf{b}$, between each input-output pair $(\tilde{\mathbf{y}}[n],\tilde{\mathbf{y}}[n+1])$ (Eq. (2.1)) in order to minimize the error between $\mathbf{Y}$ and $\tilde{\mathbf{Y}}$. To achieve the optimization we use the Lasso regression method [14], with the loss function:
@@
$\tilde{y}_{i}[n]=\sum_{j}W_{out}^{i,j}\cdot x_{j}[n]+b_{i}$ (3b)
where $\lambda$ is the lasso regularization parameter [14]; the $\langle\cdot\rangle_{n}$ notation indicates average along the time steps dimension. Then the regression objective is:
@@
where $\mathbf{w}$ denotes all the weights in matrix $\mathcal{W}_{out}$. The lasso regression method was chosen due to its effectiveness in finding the lowest required weights, preventing them from exploding, as well as effectively omitting the unnecessary weights (setting them to zero).
@@ 2.2 Retrieval of Connectivity Map
Here we demonstrate the first feature of the discussed RC-based model, which is the ability to derive the connectivity map between the measurement sites from the spatio-temporal dynamics encoded in the electrophysiological signals. Connectivity refers to the weighted relationships between the different units of the network (single neurons, populations or circuits of any type). In this analysis, we assume that this nonlinear dynamic system, can be in principle separated to linear and nonlinear regimes. Let us assume that the training of the ANN has achieved sufficiently low loss score (Eq. (2.1)), indicating that the model has learnt sufficiently well the dynamics of the given network. Looking at the dynamics described by Eq.(2.1), we note that the network state $\mathbf{y}[n]$ evolves in such way that the hypothetical connection weights between the nodes of the network are excited or depressed according to the propagation of signals within it (similarly to a biological network). We assume, however, that these complex dynamics are founded on an intrinsic network, built of fundamental connections with nominal weights. These connections could be associated with the linear regime of the dynamics. Hence, we represent Eq.(2.1), describing the network state $\mathbf{y}[n]$ evolution, by the following discrete differential equation:
$\mathbf{y}[n+1]=\mathcal{T}_{0}\mathbf{y}[n]+\mathcal{G}\Bigl{\{}\mathbf{y}[n],\mathbf{y}[n-1],\mathbf{y}[n-2]\ldots\Bigr{\}}$ (5)
where $\mathcal{T}_{0}$ is defined as the Intrinsic connectivity matrix (ICM), describing the linear relation between consecutive states. $\mathcal{G}$ denotes a nonlinear transformation that captures higher-order interactions, embodying the memory of the network. By linearizing the dynamics, it yields that:
@@
This allows to depict the morphology of this network using a directed graph of nodes and edges. Each node represents a neuronal population, whereas the edges represent the connections, with its appointed weight, where a weight of a connection represents the coupling strength between the two populations. Fig. 2 shows an example of a connectivity map obtained by the RC model and compared to the ground truth map.
In this part of the work we show the validity of the predicted ICM to represent the actual connectivity between the nodes of the network, each representing a neuronal population. We compared this prediction by the RC model with the performance of other common methods, such as Cross-Correlogram (CC) [10, 11] and Transfer Entropy (TE) [12]. As a benchmark, we built a simulation of a neuronal network with ground-truth connections between the neurons. The network has been built and simulated on NEST simulator [13], where we created virtual array of electrodes which sampled the signals from a defined population of neurons, similarly to conditions found in an experimental electrophysiological recording, such as microelectrode array (MEA). We assessed prediction accuracy through two distinct methods: 1. Receiver Operating Characteristic (ROC) Curve Analysis: This method illustrates the model’s predictive performance in a binary context, distinguishing between the presence and absence of connections. We evaluated the performance with the Area Under Curve (AUC) of the ROC curve. 2. Pearson Correlation with Ground Truth: This metric offers a quantitative measure of prediction quality by not only capturing the presence or absence of connections but also quantifying the accuracy in predicting connection weights. These weights may encompass both excitatory (positive values) and inhibitory (negative values) connections within the network.
@@
Figure 2: An example of connectivity map result obtained by the RC model. (a) A connectivity matrix and a connectivity graph map are presented for both the ground truth network (left hand side) and the connectivity map obtained by the RC model (right hand side). The Ground truth network was simulated on NEST simulator, producing electrophysiological activity which was then decoded by the RC model and obtaining the Intrinsic Connectivity Matrix (ICM). Note that the RC model predicted with a very high accuracy the existence of 3 uncoupled groups. In addition the model distinguished with a high accuracy between excitatory connections (positive weights) and inhibitory (negative weights). The nodes in the graphs represent neuronal population (or circuit) and the edges represent effective synaptic connections. The predicted map displays only weights with absolute values exceeding 0.1. (b) The corresponding ROC curve. In this example: $\rho=0.72$; $AUC=0.94$ for $\alpha=0.2$ and $m=50$.
We examined the model performance for different network configurations, such as the size of the network in terms of number of populations and the training data richness, which we defined as:
@@
$N_{t}$ is the total number of time steps used in the training and $N_{ch}$ is the number of populations equivalent to the number of channels in the experimental data. Fig. 3 summarizes the connectivity retrieval and validation results.
Figure 3: Connectivity retrieval performance. (a) RC model connectivity prediction accuracy in terms of Pearson correlation ($\rho$) and ROC AUC metrics as a function of $q$, a parameter representing the richness of the training data (Eq. (7)); (AUC: $R^{2}=0.38$), ($\rho$: $R^{2}=0.47$). (b) Comparison between the performances of different algorithms for connectivity prediction. RC model is compared to two other common methods for connectivity retrieval: Cross-Correlation (CC) and Transfer Entropy (TE). In this comparison two-sided Mann-Whitney-Wilcoxon test with Bonferroni correction was applied; AUC-RC vs. AUC-CC: $U-stat.=1820$, $p-value=4.527\times 10^{-3}$; $\rho$-RC vs. $\rho$-CC: $U-stat.=1864$, $p-value=1.568\times 10^{-3}$; AUC-RC vs. AUC-TE: $U-stat.=1993$, $p-value=4.402\times 10^{-5}$; $\rho$-RC vs. $\rho$-TE: $U-stat.=2417$, $p-value=2.534\times 10^{-12}$; (c),(d) Performance comparison between RC model, CC and TE categorized by the complexity of the network in terms of number of sampled populations, where each population represents a circuit of 5 neurons: Figure (c)- $\rho$ metric; Figure (d) AUC metric.

### 05 Mar 2024 v3

# Decoding Neuronal Networks: A Reservoir Computing Approach for Predicting Connectivity and Functionality

@@ Abstract
In this study, we address the challenge of analyzing electrophysiological measurements in neuronal networks. Our computational model, based on the Reservoir Computing Network (RCN) architecture, deciphers spatio-temporal data obtained from electrophysiological measurements of neuronal cultures. By reconstructing the network structure on a macroscopic scale, we reveal the connectivity between neuronal units. Notably, our model outperforms common methods like Cross-Correlation and Transfer-Entropy in predicting the network’s connectivity map. Furthermore, we experimentally validate its ability to forecast network responses to specific inputs, including localized optogenetic stimuli.
@@ 1 Introduction
In the field of neuroscience, electrophysiological studies offer a wide-ranging perspective on the intricate interplay among different cell types at varying scales [1]. These investigations span from understanding the function of individual cells to unraveling the dynamics within complex systems composed of numerous cells, [2], all with the aim of capturing a holistic view of brain activity. As biological systems grow in complexity, the challenge intensifies when attempting to analyze or model their behavior. To address this, a plethora of models has emerged, ranging from single-cell descriptions (such as the Hodgkin–Huxley model [3]) to large-scale population models [4, 5]. These models explore diverse aspects, including biophysical properties (e.g., membrane voltage) and information propagation (e.g., spike trains). Some innovative approaches leverage experimental data to adapt computational models that mirror biological systems [6, 7, 8]. Machine learning and deep learning techniques play a pivotal role in training these models to achieve desired outcomes. While such approaches may be computationally intensive for certain research questions, they offer a practical avenue for constructing computational tools across various applications.
@@
In this work, we propose an artificial neural network model for interpreting electrophysiological signals within intricate networks. Our approach uncovers connectivity patterns among different sampled regions in the network and captures dynamic interactions between them. The model is based on the Reservoir Computer Network (RCN) architecture [9], leveraging information gleaned from electrophysiological measurements in complex neural circuits. These circuits defy easy comprehension through standard measurement analyses, leading us to model them as nonlinear networks with internal random connections. Notably, our model outperforms common methods, including Cross-Correlation (CC) [10, 11] and Transfer-Entropy (TE) [12], in accurately predicting the network’s connectivity map (CM). Furthermore, we demonstrate the model’s ability to forecast the spatio-temporal response of a given network to specific inputs. To validate our predictions, we combine experimental measurements using microelectrode arrays (MEA) with in-vitro mice cortical neurons [13] and numerical simulations conducted using the NEST simulator [14].
2 Reservoir Computing Model for Electrophysiological Signals Decoding
@@
The Reservoir Computing (RC) architecture models electrophysiological data obtained from biological neuronal networks using multi-electrode arrays (MEA). The objective of the RC model is to decode these signals, ultimately reconstructing a network structure within the measured neuronal culture. In a MEA measurement, each electrode (channel) captures signals from a nearby neuronal population (microcircuit), as depicted in Fig. 0(a). In our artificial neural network (ANN), each such microcircuit or population is treated as a node of the macroscopic network (representing the measurement domain), and their interconnections are to be reconstructed during the training phase. This approach enables the RC model to unveil the intricate connectivity and relationships within the biological neuronal networks.
@@
To describe the ANN, let us consider a discrete time domain denoted by $n=1,2,3...$ corresponding to real time units ($t$) by the relation:
$t=n\cdot t_{int}\;.$ (1)
Since each neuronal microcircuit can be likened to a recurrent spiking neural network with a leaky integrator [5] (as shown in Fig. 0(b)), $t_{int}$ is defined as the integration time that characterizes the time duration of which a neuronal microcircuit processes incoming information. Next, we assume that the relevant node state corresponds to the firing rate measured by each MEA channel. Let us represent the network node states using a vector $\mathbf{y}[n]\in\mathbb{R}^{N_{ch}\times 1}$, where $\mathbf{y}[n]$ represents the instantaneous spike rate (ISR) measured by each of the MEA electrodes at time step $n$ (with $N_{ch}$ being the number of measurement channels). To predict the network’s next state $\mathbf{y}[n+1]$ based on the current and past states $\mathbf{y}[n],\mathbf{y}[n-1],\mathbf{y}[n-2]\ldots$, we have designed the ANN illustrated in Fig. 0(c). In this architecture, the input layer is connected to a layer of $N_{ch}$ independent reservoirs, each representing the neuronal microcircuits probed by the MEA electrodes. The reservoir has an internal dimension $m$. Additionally, each reservoir is connected to all the nodes via a trainable interconnecting matrix at the output layer. The state evolution from $\tilde{\mathbf{y}}[n]$ to $\tilde{\mathbf{y}}[n+1]$ (the tilde sign indicates predicted states) through the ANN is described by the following set of equations:
@@
  • $\mathbf{x}_{in},\mathbf{x}\in\mathbb{R}^{N_{res}\times 1}$ are the reservoir input and actual state respectively ($N_{res}=m\times N_{ch}$ is the reservoir layer dimension).
  • $\mathcal{W}_{in}\in\mathbb{R}^{N_{res}\times N_{ch}},\hat{\mathcal{S}}\in\mathbb{R}^{N_{res}\times N_{res}},\mathcal{W}_{res}\in\mathbb{R}^{N_{res}\times N_{res}},\mathcal{W}_{out}\in\mathbb{R}^{N_{ch}\times N_{res}}$ are input, synaptic, reservoir and output matrices respectively.
  • $10<\alpha<$ 1 is the reservoir layer internal memory strength.
  • $\mathbf{b}\in\mathbb{R}^{N_{ch}\times 1}$ is a bias vector.
@@
Therefore,
$\tilde{\mathbf{y}}[n+1]=\hat{\mathcal{F}}\Bigl{\{}\tilde{\mathbf{y}}[n]\Bigr{\}}\;.$ (3)
((a))
((b))
((c))
((d))
Figure 1: Description of the RC model. (a) A microscope image of amicroelectrode (diameter of 30 $\mu$m) surrounded by cultured neurons. This image illustrates how each electrophysiological measurement site samples signals from a complex circuit, whose morphology is not resolvable by simple methods. (b) An illustrative image of how the signals are intercepted by each electrode (yellow circles) from the biological neuronal network in the background. The spike trains collected at each electrode are then forming integrated signals which propagate within the network. (c) The artificial neural network (ANN) design of the RC model. It consists of three layers (input, reservoir, output) and a recurrent branch. It describes how the state of the network $\mathbf{y}$ (described by $N_{ch}$ samples of the measured signals from each electrode) is processed from time step $n$ to time step $n+1$. (d) A block diagram illustrating the sequential computational processes, including training, validation, and testing, conducted in this study.
To train the RC model, we utilized both experimental and simulated ISR sequences. Given that the ANN belongs to the RCN type, training is exclusively applied to the output layer (Eq. (2c)) using a linear regression. This leverages the intricate dynamics embedded in the reservoir layer (described by Eqs. (2a),(2b)). Let us consider the training data matrix $\mathbf{Y}\in\mathbb{R}^{N_{ch}\times N_{t}}$, which governs $N_{t}$ time steps of the observed network state. In other words, $\mathbf{Y}=\Bigl{[}\mathbf{y}[1],\mathbf{y}[2]\ldots\mathbf{y}[N_{t}]\Bigr{]}$. The training task involves optimizing the matrix of weights $\mathcal{W}_{out}$ and the biases $\mathbf{b}$ of the output layer, in order to minimize the error between $\mathbf{Y}$ and $\tilde{\mathbf{Y}}$. We use the Lasso regression method [15], employing the following loss function:
@@
$\tilde{y}_{i}[n]=\sum_{j}W_{out}^{i,j}\cdot x_{j}[n]+b_{i}\;.$ (4b)
Here $\lambda$ represents the lasso regularization parameter, and the $\langle\cdot\rangle_{n}$ notation denotes averaging along the time steps dimension. Then the objective of the regression is expressed as:
@@
where $\mathbf{w}$ encompasses all the weights in the matrix $\mathcal{W}_{out}$. The choice of the lasso regression method stems from its effectiveness in identifying essential weights (preventing them from exploding) and efficiently excluding unnecessary weights by setting them to zero.
3 Results
@@ 3.1 Retrieval of Connectivity Map
Here, we demonstrate the first feature of our Reservoir Computing (RC) model: the ability to derive the connectivity map between measurement sites from the spatio-temporal dynamics encoded in electrophysiological signals. Connectivity refers to the weighted relationships between different nodes of the network.
Let us assume that the Artificial Neural Network (ANN) has successfully learned the dynamics of a given biological network (i.e., achieved low loss values during training). The dynamics described by Eq.(2) show the evolution of the network state $\mathbf{y}[n]$, which occurs on an intrinsic network composed of fundamental connections with varying strengths. We hypothesize that these connections can be associated with the linear regime of the dynamics. Therefore, we represent the nonlinear dynamics described by Eq.(3), which governs the evolution of the network state $\mathbf{y}[n]$, as follows:
$\mathbf{y}[n+1]=\hat{\mathcal{F}}\Bigl{\{}\mathbf{y}[n]\Bigr{\}}=\mathcal{T}_{0}\mathbf{y}[n]+\mathcal{G}\Bigl{\{}\mathbf{y}[n],\mathbf{y}[n-1],\mathbf{y}[n-2]\ldots\Bigr{\}}$ (6)
where $\mathcal{T}_{0}$ is defined as the Intrinsic connectivity matrix (ICM), describing the linear relation between consecutive states. $\mathcal{G}$ denotes a nonlinear transformation that captures higher-order interactions, embodying the memory of the network. By linearizing Eq.(2), we obtain:
@@
This allows us to depict the morphology of the network using a directed graph of nodes and edges. Each node represents a neuronal population, whereas the edges represent the connections, with their appointed weights. The weight of a connection signifies the coupling strength between the two populations.
((a))
((b))
((c))
@@
Figure 2: Connectivity map obtained by the RC model trained on NEST simulation data. (a), (b) Connectivity matrix (left) and connectivity graph map (right) for both the ground-truth network (a) and the the RC model (b). Color bar refers to the connection strengths (weights). The ground-truth network was simulated by a NEST simulator, producing an electrophysiological activity which was then decoded by the RC model obtaining the Intrinsic Connectivity Matrix (ICM). The nodes in the graphs represent neuronal population (or circuit) as probed by the electrodes (channels) and the edges represent their connections. Connection strengths in the RC model are deduced by the weight matrix $\mathcal{T}_{0}$ (Eq. 7). Each weight was normalized by the highest element in absolute value and color coded between -1 and 1. Note that the RC model predicted with a very high accuracy the existence of 3 uncoupled groups of nodes. In addition, the model distinguished with a high accuracy between excitatory connections (positive weights) and inhibitory (negative weights) yielding a Person correlation $\rho=0.72$. The connectivity map in (b) displays only weights with absolute values exceeding 0.2. (c) ROC curve to which corresponds a $AUC=0.94$. In this example, the RC model had a memory strength $\alpha=0.5$ and a reservoir dimension $m=50$.
To validate our hypothesis, we utilize the ICM to represent the actual connectivity between the nodes of the network, with each node corresponding to a neuronal population. As a benchmark, we construct a simulation of a neuronal network with ground-truth connections between groups of neurons (named ground-truth network). This network is built and simulated using the NEST simulator [14]. In the simulation, we create a virtual array of electrodes that sample signals from a defined population of neurons. This setup mirrors conditions found in experimental electrophysiological MEA recordings. Next, we benchmark the RC model with both the ground-truth network and experimental data. We evaluate prediction accuracy using two distinct methods:
  1. 1.
    Receiver Operating Characteristic (ROC) Curve Analysis [16, 17]. This method illustrates the model’s predictive performance in a binary context, distinguishing between the presence and absence of connections among the nodes. We assess the performance using the Area Under the Curve (AUC) of the ROC curve, where a value of 1 means that all the connections are properly detected by the RC model.
  2. 2.
    Pearson Correlation ($\rho$) with ground-truth [18, 19]. This metric provides a quantitative measure of the model’s prediction quality. It not only captures the presence or absence of connections among the nodes but also quantifies the accuracy in predicting connection weights. These weights may encompass both excitatory (positive values) and inhibitory (negative values) connections within the network.
To effectively elucidate the model’s capacity, we showcase its performance by simulating the MEA signals from a neuronal population formed by three distinct, non-interconnected groups of nodes (node=neurons population probed by a single electrode). Each group of 20 nodes (out of a total of 60) possesses internal connections, while inter-group connections are absent. Figure 1(a) shows the input connectivity matrix or connectivity graph map used to simulate the data used to train the RC model. After training, by linearizing the RC model we extracted the the connectivity map shown in Fig. 1(b) which is compared to the ground-truth maps (Fig. 1(a)). The obtained accuracy results of $AUC=0.94$ (Fig. 1(c)) and $\rho=0.72$ (Pearson correlation) demonstrate the capacity of the RC model in predicting the connectivity. It not only distinguishes the existence and absence of connections ($AUC$) but also retrieves the weight values and distinguishes between excitatory and inhibitory connections (positive or negative sign) ($\rho$). In this way we can use MEA recordings to decipher intricate connectivity patterns by the analysis of spatio-temporal signals from an actual neuronal culture. An example of a connectivity map obtained by the RC model from experimental MEA data can be found in Fig. 3.
Figure 3: Connectivity map derived by the RC model from experimental data collected using a 60-electrode MEA and a mice cortical neuronal culture. The connectivity weights are normalized based on the highest absolute value within the obtained ICM model. The color code refers to the color bar on the right, scaled between -1 and 1. To enhance clarity, connections with weights below $|0.25|$ have been omitted.
Furthermore, we compared the RCN model’s performance with the Cross-Correlogram (CC) [10, 11] and Transfer Entropy (TE) [12] methods. These methods are often used to analyze neuronal interactions and connectivity patterns. Our investigation spanned various network configurations, (a total of 51 simulations), including different network sizes, morphologies (sparse vs. dense populations), and ratios of inhibitory to excitatory connections. These structural variations led to significant differences in electrophysiological dynamics, such as spike and burst rates. To gauge the RC model’s effectiveness, we introduced a parameter $q$, which reflects the richness of the training data set:
@@
where, $N_{t}$ is the total number of time steps used in the training of the RC model and $N_{ch}$ is the number of population clusters (nodes) in the ground-truth network or the number of MEA channels in the experimental data. Fig. 4 summarizes the connectivity retrieval and validation outcomes. In Fig. 3(a), we observed a notable correlation between the connectivity retrieval performance and the parameter $q$. This suggests that for larger networks, more time steps in the dynamics need to be included during training. Additionally, the predictive ability of the RC model is influenced by the sparseness of neuronal populations. Figure 3(b) clearly demonstrates that the RC model consistently outperforms the CC and TE methods in terms of connectivity retrieval. Further insights are provided by Figs. 3(c) and 3(d). These figures break down algorithm performance with respect to network size, considering $\rho$ and $AUC$, respectively. Notably, as the number of populations increases, the competing algorithms (CC and TE) exhibit a decline in performance. In contrast, the RC model remains less sensitive to network size, indicating its robustness across varying scales.
((a))
((b))
((c))
((d))
Figure 4: Connectivity retrieval performance of the various algorithms with respect to the ground-truth network. (a) Pearson correlation ($\rho$) and ROC AUC metrics as a function of the richness training data set parameter $q$ (Eq. (8)) for the RC model. Points refer to various network realizations, the lines are a fit to the data (AUC: $R^{2}=0.38$ and $\rho$: $R^{2}=0.47$), the bands refer to the 95 % confidence interval. (b) Comparison between the performances of the different algorithms evaluated on all the 51 different network realizations. RC refers to the RC model, CC to the Cross-Correlation method and TE to the Transfer Entropy method. Box plots are used to visualize the distribution of the data (minimum value, 1st quartile, median, 3rd quartile, and maximum value). Diamonds are out-layer points. To gauge the statistical relevance of the algorithm comparison, two-sided Mann-Whitney-Wilcoxon test with Bonferroni correction was applied; AUC-RC vs. AUC-CC: $U-stat.=1820$, $p-value=4.527\times 10^{-3}$; $\rho$-RC vs. $\rho$-CC: $U-stat.=1864$, $p-value=1.568\times 10^{-3}$; AUC-RC vs. AUC-TE: $U-stat.=1993$, $p-value=4.402\times 10^{-5}$; $\rho$-RC vs. $\rho$-TE: $U-stat.=2417$, $p-value=2.534\times 10^{-12}$; (c),(d) Comparison between the performances of the different algorithms as a function of the number of populations (network nodes), where each population represents a circuit of 5 neurons.